FEL Upgrade Project Weekly Brief
January 7-11, 2002
Happy New Year to all of our colleagues. Within the intense pace of activities for the FEL Upgrade construction project. we made final preparations for next week's FEL Users/Laser Processing Consortium Workshop on Jan. 17-18. The full program for the workshop is given on http://www.jlab.org/FEL/LPC/lpcagenda02jan.html. All interested parties are invited to this annual event.

All three 100 kW klystrons were successfully tested to 50 % power levels (the limits of our presently installed power).

Discussions were held this week with our JTO and ONR contract monitors for documentation needed for FY02 support for presently unfunded FEL upgrade activities including: (1) final system commissioning; (2) completion of the third cryomodule; and (3) modification or addition of a new wiggler for optimized high power operation at 1- 2 microns.

The project monthly technical and financial reports for November 2001 were completed and forwarded to the DOE, ONR and AFRL contract monitors.

The tasking statements for the newly awarded JTO proposals on cryogenic mirror development

and drive laser development were received from JTO and returned with comments so that funding for these tasks can be added to our existing or a revised ONR/DOE Memorandum of Agreement.

Michelle Shinn and George Neil made presentations on our work on laser materials damage studies at the JTO Project Continuation Review this week (Jan. 8-9) held in Albuquerque.

WBS 4 (Injector):
The gun chamber had its top and bottom flanges welded on and the spider locating groove is being machined. Tolerances on the gun stand drawings were loosened up and the package is back out for fabrication bids. The gun mounting ring was submitted to the shop for fabrication. W&M is running a test implant on a 3" diameter tube prior to implant of the cathode support tube. Solenoid mechanical modifications are being designed to facilitate placing correctors inside of the solenoid.

WBS 6 (RF):
Quarter Klystrons - All 3 of the 100 kW klystrons were successfully tested at half power with the existing HVPS. Modifications are being made to the klystron cart design to make the piping more compact. The LCW headers in the RF gallery are being drained today so larger taps can be installed for the klystrons, circulators, RF loads, and HOM loads.

Quarter HVPS - The rigging contractor on next Wednesday plans to remove the existing transformers, inductors, and variacs and install the new transformers (3000 lbs) and inductors (1600 lbs).

Master Oscillator - The Master Oscillator is to be upgraded (more output power) and extended to Zones 3 & 4.

2G Dump - With RadCon's permission the water bottles were removed, drained, and stored in their plastic milk cartons. This was necessary before the 2G Dump can be dismantled. This dump is still activated and its removal will require RadCon's supervision.

WBS 8 (Instrumentation):
Several manhours have been spent this week removing and re-routing cables in preparation for cable tray modifications to support the new machine.

Six beamviewers have been assembled and delivered to the vacuum group for mounting on their respective pedestals.

The new inventory control and barcoding data base is nearly functional. There are some security items that need to be added before it becomes available for general access.

Design of a new 8 channel Stepper Motor Control Chassis is complete and should be ready to start fabrication next week.

WBS 9 (Transport):
Optical Chicane Dipoles (GW)
o At Magnet Enterprises International in Oakland CA., contrary to what was reported earlier, the potting of the first GW coil had some dry spots. They are making a new coil and redoing their potting mold to eliminate the problems. The potting of the new coil should begin next week. Machining of cores and parts has started.
Injector Dipoles (DU/DV)
o Procurement is sending out the bid for the magnets today. It will be an electronic procurement as was successful with the 180 degree Dipole (GY).
Arc Dipoles (GY, GX, GQ)
o Dave Douglas accepted the magnetic design of the Reverse Bends (GQ). He has asked Advanced Energy Systems (AES) to give him more field integral information on the switch mode (to the UV Line) of the Bend Magnet (GX) before he accepts this design. This variation of the powering uses only half of each coil at full current to reduce the field for the electron beam to angle into the UV beam line. The vellums of all the drawings have been prepared for signature and the specifications are in composition.
o Steel was shipped and received for the 180° Dipoles (DY) to Bosma Inc. of Ohio for manufacturing the Cores. Copper was shipped Wang NMR of CA for coil manufacturing.
Magnet Measurement
o Measurement of the dipole chamber's magnetic properties started up using weld samples prepared by the company that welded the 180° chambers last time.
o Design is nearly complete for the probe system to measure absolute field of the Bend and Reverse Bend Dipoles.
3 inch quad (QX)
o The core and assembly vender, New England Technicoil (NTC) shipped the first 10 completed quadrupoles.
o Magnet Enterprises International (MEI) shipped 60 more coils to New England Tecnichoil.
Trim Quad (QT)
o The magnets continue in fabrication at Milhous Control of Virginia.
Sextupole (SF)
o DULY Research has a magnetic model for the sextupole that achieves specification. They are now working at superposing the vertical dipole corrector field onto the magnet.
Octupole (OT)
o Work on this magnet is on furlough until the sextupole is designed.
Corrector Dipoles (DB, DJ)
o These magnets continue to be qualified on the measurement stand.
Beam Line and Vacuum
o Work is going very well at AES on the design the Arc Chambers.
o Design work on girders and stands continued. Parts for the regions between the cryomodules are finally in checking.
o The girders for return leg quadrupoles are being assembled in the new space available in the addition to the Test Lab.

WBS 11 (Optics):
We received the optical cavity stepper motors with integral PZTs before the end of the year. We’re building a small enclosure so we can preserve the UHV rating and begin testing. We also received ball bearings treated by Dicronite, and we will begin testing them for particulate generation. Review of the drawing package for the optical cavity assemblies revealed a few parts that, with small changes, would yield items that are simpler to fabricate. The revision of the ultraviewer mirror mount has started. We moved the OCCMS work to the User Lab 6 hutch, where we can take advantage of the enclosure to dampen out air currents.

We have devoted some time to eliminating air bubbles in the glue joint for the deformable mirror. By degassing the mixture, we have been able to drop the density dramatically (from  ~100 to 3/sq. in.).

Over the holidays we held e-mail meetings with a possible vendor for the drive laser oscillator.

Operations/Decommissioning/Other Activities:
We received a power meter from Molectron that had been sent out for repair and recalibration.